Determine whether the map T : R² + R² defined by ſa + 367 T ((o) a – b is a linear transformation.
Answer #1

a +38 a-b = V= 4-6 + 4x +38] Give the map 7: R2 R2 defined by 7 we know the linear transformation: T(u+v)=7(u)+7(V) T(cu)=cT(